OpenStudy (turingtest):

Prove 1=0.999... Do not answer if you have seen this question before! Figure it out!

OpenStudy (anonymous):

1/3 = 0.33333... 1 = 1/3 x 3 0.33333... x 3 = 0.99999.... 1 = 0.9999....

OpenStudy (turingtest):

Very nice! The way I know to do it is let x=0.999... 10x=9.9999 9x=9 x=1 pretty much equivalent. You're good, what do you study?

OpenStudy (anonymous):

I am still in high school

OpenStudy (turingtest):

damn, and you figured this out on your own?

OpenStudy (turingtest):

you didn't cheat by looking online did you?

OpenStudy (anonymous):

nope, this is kind of basic comparing to the math contests I do in school...

OpenStudy (turingtest):

Yeah I used to do those non-standard tests for this kind of thing in high school too. Still, very nice. I'm sure you'll be a kick-retricemathematician.

OpenStudy (turingtest):

kick-retrice lol where did "retrice" come from?

OpenStudy (turingtest):

oh it won't let me type "A S S"

OpenStudy (anonymous):

lol I understand. Thank you

OpenStudy (jamesj):

@moneybird, are you in Australia?

OpenStudy (anonymous):

No, canada

OpenStudy (jamesj):

got it. I thought your koala might mean you were.

OpenStudy (agreene):

Pretty sure every proof I've seen for this is here: http://en.wikipedia.org/wiki/0.999...